Hrubec Products, Inc., operates a Pulp Division that manufactures wood pulp for use in the production of various paper goods. Revenue and costs associated with a ton of pulp follow:

Selling price $ 25
Expenses:
Variable $ 16
Fixed (based on a capacity of
97,000 tons per year)
6 22
Net operating income $ 3

Hrubec Products has just acquired a small company that manufactures paper cartons. This company will be treated as a division of Hrubec with full profit responsibility. The newly formed Carton Division is currently purchasing 31,000 tons of pulp per year from a supplier at a cost of $25 per ton, less a 10% purchase discount. Hrubec’s president is anxious for the Carton Division to begin purchasing its pulp from the Pulp Division if an acceptable transfer price can be worked out.

Required:

For (1) and (2) below, assume the Pulp Division can sell all of its pulp to outside customers for $25 per ton.

1. What is the lowest acceptable transfer price from the perspective of the Pulp Division? What is the highest acceptable transfer price from the perspective of the Carton Division? What is the range of acceptable transfer prices (if any) between the two divisions? Are the managers of the Carton and Pulp Divisions likely to voluntarily agree to a transfer price for 31,000 tons of pulp next year?

2. If the Pulp Division meets the price that the Carton Division is currently paying to its supplier and sells 31,000 tons of pulp to the Carton Division each year, what will be the effect on the profits of the Pulp Division, the Carton Division, and the company as a whole?

For (3)–(6) below, assume that the Pulp Division is currently selling only 56,000 tons of pulp each year to outside customers at the stated $25 price.

3. What is the lowest acceptable transfer price from the perspective of the Pulp Division? What is the highest acceptable transfer price from the perspective of the Carton Division? What is the range of acceptable transfer prices (if any) between the two divisions? Are the managers of the Carton and Pulp Divisions likely to voluntarily agree to a transfer price for 31,000 tons of pulp next year?

4-a. Suppose the Carton Division’s outside supplier drops its price (net of the purchase discount) to only $21 per ton. Should the Pulp Division meet this price?

4-b. If the Pulp Division does not meet the $21 price, what will be the effect on the profits of the company as a whole?

5. Refer to (4) above. If the Pulp Division refuses to meet the $21 price, should the Carton Division be required to purchase from the Pulp Division at a higher price for the good of the company as a whole?

6. Refer to (4) above. Assume that due to inflexible management policies, the Carton Division is required to purchase 31,000 tons of pulp each year from the Pulp Division at $25 per ton. What will be the effect on the profits of the company as a whole?
Answer #1

Lowest Acceptable transfer price for Pulp Division = $25 i.e.Selling Price to outside customers

Highest Acceptable transfer price for Carton Division = $25*0.9 = $22.5 (current purchase price)

No Acceptable Range

Not likely to Agree

2.Effect on pulp Division:

Loss on 31,000 tons = (25-22.5)*31,000 = $77,500

Profits of pulp division will decrease by $77,500

Carton Division = No effect (same price as before)

Company as a whole = Loss of $77,500

i.e. profits will decrease by $77,500

3.Space Capacity Exists

Lowest Acceptable transfer price = Variable Cost = $16

Highest Acceptable Transfer Price for Carton Division = $25*0.9 = $22.5 (current purchase price)

Range of acceptable transfer price = $16-$22.5

Yes, they are likely to voluntarily agree

4a – The price is still higher than the variable cost to the pulp division, it should meet this price.

4b Effect on profits of the company is equal to loss of $5 per ton

=31,000*$5 = loss of $155,000

i.e. profits will reduce by $155,000

5. The carton division can accept a higher price for the good of the company as a whole

6. Effect on profits of the company:

Increase in profits of pulp division = 31,000*(25-16) = $279,000

Loss of Carton Division = 31,000*(25-21) = $124,000

Net benefit to the company = $155,000

Hence, profits will increase by $155,000